# Towards Understanding Link Predictor Generalizability Under Distribution Shifts

### v2

@@ 5.2. Results for GNN4LP
Observation 2: Performance Differs By Split Type and Threshold. As shown in Figure 4, regardless of whether a model is tested on a “Forward” or “Backward” split of ogbl-ppa; the change in structural information for each subsequent split gradually changes a model’s performance. On the “CN” split, a stark increase is seen between ”Split 1” and ”Split 2”, indicating that more structural information between training and validation improves GNN4LP performance (Wang et al., 2023a), even on dense datasets like ogbl-ppa. The fact that these results include splits produced by Preferential-Attachment, Global Structural Information (SP), and Local Structural Information (CN) indicates the effect of any change in structural information when training LP models (Mao et al., 2024b).
Observation 3: Performance Impact Varies by Model. All raw MRR scores are stored in Appendix Section H within Tables 7 to 9. We can make the following observations by split type. Common Neighbors: Most models fail to generalize on the “Backward” CN splits. However, once more Common Neighbors are made available in the CN - 4,2 and CN - 5,3 splits; NCNC performs 2 to 3 times better than other GNN4LP models. Therefore, indicating that it is possible to generalize with limited local information. Shortest-Path: GNN4LP Models which rely more on local structural information (i.e. NCNC, LPFormer, and SEAL) typically suffer more under the “Backward” SP splits, resulting in the models performing 2x to 4x worse than BUDDY or GCN. Therefore, indicating the necessity for models to adapt in scenarios with an absence of local structural information. Preferential-Attachment: LPFormer performance on the PA split is 2 times higher than HeaRT’s ogbl-collab (Li et al., 2024), but reduces drastically on LPShift’s ogbl-ppa or ogbl-ddi. Therefore, indicating the impact that structural shift incurs with the increasing density of datasets.
@@ 5.3. Results for Generalization Methods
Observation 1: Graph-Specific Generalization Methods Can Help. As demonstrated in Table 2, the two generalization methods specific to LP: TC (Wang et al., 2023b) and EPS (Singh et al., 2021) fail to increase performance under LPShift in the ”Forward” CN and PA splits. However, the methods do improve performance, as indicated by the ”Backward” DDI split, where both EdgeDrop and EPS increase model performance by 3.87% and 7.5%, respectively. This is likely due to the exceptionally dense nature of DDI, where numerous high-degree nodes are likely made more relevant to performance through: 1) the removal of edges from EdgeDrop or 2) connecting relevant neighbors with EPS. To validate this, we calculate Earth Mover’s Distance (EMD) (Rubner et al., 1998) between the heuristic scores of the training and testing splits before and after applying the generalization methods. EPS injects CNs into the training adjacency matrix, significantly altering the training and testing distributions. This drastic change is indicated in Figures 8 and 5 with the near-perfect EMD scores of EPS. Such a change in EMD and the Table 2 results, demonstrates that generalizing under LPShift can be achieved with targeted augmentations to the graph data structure. However, EdgeDrop has relatively-minimal effect on the EMD scores, indicating that non-targeted augmentations can function on highly-dense datasets.
[…]
@@
Observation 2: Traditional Generalization Methods Work in Limited Scenarios. As shown in Table 3, any of the generalization methods tested on the ”Forward’ split rarely improve performance. IRM and DRO oftern reduce performance the most, sometimes by more than 10% MRR. All methods encounter more consistent performance improvements on the ”Backward” ogbl-ppa and ogbl-ddi splits. Given that all generalization methods are trained on environmental subsets defined by structural heuristics, it follows that the traditional generalization methods improve performance under LPShift when given sample subsets with distinct and diverse structural patterns. Future work could explore this idea further by applying different heuristic scores to each environmental subset.
@@
Figure 6. Three subplots corresponding to: 1.) the ’CN - 1,2’ LPShift on ogbl-collab 2.) the ’CN - 2,1’ LPShift on ogbl-collab 3.) CN predictor performance for the original ogbl-collab, ’CN - 1,2’, and ’CN - 2,1’ splits.
@@
Three subplots corresponding to: 1.) the ’CN - 1,2’ LPShift on ogbl-collab 2.) the ’CN - 2,1’ LPShift on ogbl-collab 3.) CN predictor performance for the orginal ogbl-collab, ’CN - 1,2’, and ’CN - 2,1’ splits.
5.4. Discussion
Does GNN4LP generalize and do generalization methods work? As detailed in Table 1, LPShift significantly reduces the performance for a majority of LP methods versus HeaRT. This is especially notable given the difficulty HeaRT imposes on the original benchmark setting (Hu et al., 2020). This observation coupled with the limited success of all generalization methods on ”Forward” splits in Tables 2 and 3 shows the difficulty posed on LP models when trained with limited structural information relative to evaluation samples.
How is the proposed distribution shift affecting performance? In order to verify that LPShift’s effect on dataset structure is significant and can be correlated to performance, we follow a similar analysis conducted in (Wang et al., 2023a), as shown in Figure 6. In which, CN’s predictive performance is measured and compared under the ”Forward” and ”Backward” LPShift. We choose CN as the predictor given that it relies on directly ranking the given dataset samples by structure in order to predict if a link forms; meaning performance is sensitive to a change in dataset structure.
The first two subplots in Figure 6 depict the density estimates of the CN distribution for the CN - 1,2 and CN - 2,1 LPShift splits of the ogbl-collab dataset. The third subplot depicts the performance of CN on the datasets depicted in the first two subplots.
Given the targeted effect of LPShift on dataset structure, the reduced performance (1% MRR) of the CN ranking, as shown in Figure 6, on the CN - 1,2 split is due to the lack of structural information in the training split (CN = 0), limiting it’s ability to rank valid and test samples with $>$1 CNs. The structure-to-performance link is clearer in the CN - 2,1 scenario: where training with 2+ CNs yields 60% MRR, but fails to generalize on test samples with zero CNs, shown by the 1% MRR score. Given that all GNN4LP models incorporate neighborhood/structural information into their architectures (Wang et al., 2023a), it follows that GNN4LP’s reduced performance under LPShift is due to the splitting strategy enforcing a scenario where models must generalize to dataset samples drawn from a different distributions.
@@ 6. Conclusion
This work proposes LPShift, a simple dataset splitting strategy for inducing structural shift relevant for link prediction. The effect of this structural shift was then benchmarked on 16 shifted versions of ogbl-collab, ppa, and ddi, posing a unique challenge for SOTA GNN4LP models and generalization methods versus the HeaRT. Further analysis indicates that current generalization methods improve performance under LPShift in scenarios with diverse pairwise structures amongst training samples. As such, LPShift demonstrates that GNN4LP is vulnerable to distribution shift induced by any pairwise heuristic. Given the severity of structural shift induced by LPShift, future work can consider relaxing the threshold between sample distributions or consider mixing structural heuristics to better mimic real-world distribution shift and the complexities therein.
Appendix A Realistic Scenarios for Structural Shift in Link Prediction
  • Adversarial Recommender Systems = Shortest-Path: A company may want to understand which products to avoid showing to a potential customer without need to hear the user’s preferences directly. In this scenario, global information, as captured by Shortest-Path, becomes the most valuable for the specific use-case.
  • Social Recommender Systems = Preferential-Attachment: A video-streaming platform working with independent content creators may wish to understand what drives users to engage with the platform’s content creators, so that engagement can increase for less-popular creators. As a starting point, the content creators with the most followers may have different characteristics that increase engagement versus less-popular creators. So, the streaming platform may wish to tune their dataset to determine if their recommendation system can generalize from more to less-popular creators.
  • Movie Recommender Systems = Common-Neighbors: A movie-streaming platform wants to provide suggestions to users that are the most relevant to the user’s current interests. So, the movie platform sorts possible movie recommendations by how much overlap the movies share with one another. In order to enhance exposure to new movies that overlap with a user’s interests, the streaming platform can apply the Common Neighbor LPShift to their dataset to force the algorithm to generalize in a scenario where movies may not fully-overlap.
Appendix B Heuristic Choice
Resource Allocation and Adamic-Adar Index were not considered for splitting strategies given that they build upon the original Common Neighbor formulation. Their inclusion is redundant given our intentions to induce distinctive structural shifts based on varying structural information, as described for each heuristic in Section 3.
Appendix C Splitting Strategy – Additional Algorithmic Details
This section provides additional details about the way data was formatted before being used as input for Algorithm 1 of our proposed splitting strategy and the intuition behind how Preferential-Attachment and Shortest-Path work within the splitting strategy. The details on the algorithm includes:
  • Validation and Testing Edges are limited to 100k edges total.
@@ Appendix C Splitting Strategy – Additional Algorithmic Details
  • Negative Testing and Validation edges are produced via HeaRT (Li et al., 2024).
@@
Common Neighbors, Preferential-Attachment and Shortest-Path, as shown in Figure 2(a), 2(b), and 2(c) respectively, are interchangeable within the dataset splitting strategy. Details about how Common Neighbors functions within the strategy are included in Section 3.2. Figure 2(b) and Figure 2(c) serve as toy examples and do not correspond directly to any dataset splits tested within our study. However, the examples illustrated within Figure 2(b) and Figure 2(c) do correspond to how their given heuristic functions within our splitting strategy.
@@
For Figure 2(b) or Preferential-Attachment, it determines the degrees between a given source and target node and then multiples the two to produce the score, based on that score, the sample is then sorted into a new dataset split.
@@
For Figure 2(c) or Shortest-Path, the heuristic determines the score by determining the minimum number of nodes necessary to reach the target node from the source node. If there is a link between the two nodes, we remove the link and then re-add to the adjacency matrix after the score calculation. The final Shortest-Path score applies the calculated shortest-path length, $SP(u,v)$ as the denominator in a ratio of $\frac{1}{SP(u,v)}$, which is then used to sort the sample into it’s respective dataset split.
Appendix D Additional Analysis Details
Note: We were unable to test CFLP (Zhao et al., 2022) and EERM (Wu et al., 2022) after adapting their current implementations to our evaluation settings. CFLP experienced an out-of-memory error on all tested dataset splits. EERM experienced an out-of-memory error on every LPShift split of ogbl-ppa and exceeded 48 hours per run on ogbl-collab before converging.

### v3

@@ 5.2. Results for GNN4LP
Observation 3: Performance Impact Varies by Model. All raw MRR scores are stored in Appendix F within  Tables 9 to 11. From those tables, we can make the following observations by split type. Common Neighbors: Most models fail to generalize on the “Backward” CN splits. However, once more Common Neighbors are made available in the CN - 4,2 and CN - 5,3 splits; NCNC performs 2 to 3 times better than other GNN4LP models. Therefore, indicating that it is possible to generalize with limited local information. Shortest-Path: GNN4LP Models which rely more on local structural information (i.e. NCNC, LPFormer, and SEAL) typically suffer more under the “Backward” SP splits, resulting in the models performing 2x to 4x worse than BUDDY or GCN. Therefore, indicating the necessity for models to adapt in scenarios with an absence of local structural information. Preferential-Attachment: LPFormer performance on the PA split is 2 times higher than HeaRT’s ogbl-collab (Li et al., 2024), but reduces drastically on LPShift’s ogbl-ppa or ogbl-ddi. Therefore, indicating the impact that structural shift incurs with the increasing density of datasets. These observations within Section 5.2 provide implications for RQ2, showing that targeted model updates; like NCNC applying completion with CNs (Wang et al., 2023a) or LPFormer’s PPR-aware attention (Shomer et al., 2024; Brin and Page, 1998) aid generalization within CN and PA distribution shifts.
@@ 5.3. Results for Generalization Methods
Observation 1: Graph-Specific Generalization Methods Can Help. As demonstrated in Table 2, the two generalization methods specific to LP: TC (Wang et al., 2023b) and EPS (Singh et al., 2021) fail to increase performance under LPShift in the ”Forward” CN and PA splits. However, the methods do improve performance for the ”Backward” DDI split, where both EdgeDrop and EPS increase model performance by 3.87% and 7.5%, respectively. This is likely due to the exceptionally dense nature of DDI, where numerous high-degree nodes are likely made more relevant to performance through: 1) the removal of edges from EdgeDrop or 2) connecting relevant neighbors with EPS. To validate this, we calculate Earth Mover’s Distance (EMD) (Rubner et al., 1998) between the heuristic scores of the training and testing splits before and after applying the generalization methods. EPS injects CNs into the training adjacency matrix, significantly altering the training and testing distributions. This drastic change is indicated in Figures 5 with the 28-92% improvement of EMD scores between training and testing from applying EPS. Such a change in EMD, coupled with the Table 2 results, demonstrates that generalizing under LPShift can be achieved with targeted augmentations to the graph data structure. However, EdgeDrop has relatively-minimal effect on the EMD scores, indicating that non-targeted augmentations can generalize from highly-dense training subsets to less-dense test subsets.
[…]
Observation 2: Traditional Generalization Methods Work in Limited Scenarios. As shown in Table 3, any of the generalization methods tested on the ”Forward’ split rarely improve performance. IRM and DRO often reduce performance the most, sometimes by more than 10% MRR. All methods encounter more consistent performance improvements on the ”Backward” ogbl-ppa and ogbl-ddi splits. Given that all generalization methods are trained on environmental subsets defined by structural heuristics, it follows that the traditional generalization methods improve performance under LPShift when given sample subsets with distinct and diverse structural patterns. Future work could explore this idea further, applying different heuristic scores to each environmental subset or extending invariant learning principles to work for the link prediction task (Wu et al., 2024). As shown in Table 3, IRM performs well on ”Backward” ogbl-ddi splits but fails on ”Backward” ogbl-collab splits. Indeed, there is no single generalization method which functions under all LPShift scenarios for each given dataset. This result is consistent with the application of OOD generalization techniques in other data domains (Gulrajani and Lopez-Paz, 2020). As such, this result also answers RQ3 by showing LPShift poses a challenge to current generalization methods.
@@
5.4. Discussion
Does GNN4LP generalize and do generalization methods work? As detailed in Table 1, LPShift significantly reduces the performance for a majority of GNN4LP models versus the HeaRT baseline. This is especially notable given the difficulty HeaRT imposes on the original benchmark setting (Hu et al., 2020). This observation, coupled with the limited success of all generalization methods on ”Forward” splits in Tables 2 and 3, shows the difficulty posed on LP models when trained with limited structural information relative to evaluation samples.
How is the proposed distribution shift affecting performance? In order to answer RQ4 and verify that LPShift’s effect on dataset structure is significantly correlated to performance, Figure 6 follows a similar analysis conducted in (Wang et al., 2023a). In which, CN’s predictive performance is measured and compared under the ”Forward” and ”Backward” LPShift. We choose CN as the predictor, given that directly ranks given dataset samples by structure to predict if a link forms; meaning it’s predictive performance is sensitive to a change in dataset structure.
Figure 6. Three subplots corresponding to: Top.) CN predictor performance for the original ogbl-collab with it’s respective LPShift ’CN - 1,2’, and ’CN - 2,1’ splits., Bottom-Left.) the density of Common Neighbors for ’CN - 1,2’ LPShift on ogbl-collab, Bottom-Right.) the density of Common Neighbors for ’CN - 2,1’ LPShift on ogbl-collab.
@@ 5.4. Discussion
The bottom two subplots in Figure 6 depict the density estimates for CN distributions of LPShift’s CN - 1,2 and CN - 2,1 splits on the ogbl-collab dataset. The topmost subplot depicts the performance of CN on the datasets depicted in the bottom two subplots, as well as the original ogbl-collab dataset.
Figure 6 demonstrates the targeted effect of LPShift on dataset structure, the reduced performance (1% MRR) of the CN ranking on the CN - 1,2 split is due to the lack of structural information in the training split (CN = 0), limiting CN’s ability to rank valid and test samples with $>$1 CNs. The structure-to-performance link is clearer in the CN - 2,1 scenario: where training with 2+ CNs yields 60% MRR, but fails to generalize on test samples with zero CNs, shown by the 1% MRR score. Given that all GNN4LP models incorporate neighborhood/structural information into their architectures (Wang et al., 2023a; Chamberlain et al., 2022), it follows that GNN4LP’s reduced performance under LPShift is due to the splitting strategy enforcing a scenario where models must generalize to dataset samples drawn from a different distributions, empirically verifying Theorem 4.2.
@@ 6. Conclusion
This work proposes LPShift, a simple dataset splitting strategy for inducing structural shift relevant for link prediction. The effect of this structural shift was then benchmarked on 16 shifted versions of ogbl-collab, ppa, and ddi, posing a unique challenge for SOTA GNN4LP models and generalization methods versus the HeaRT baseline. Further analysis indicates that current generalization methods improve performance under LPShift in scenarios with diverse pairwise structures amongst training samples. This indicates promising future directions for targeted graph data augmentation and invariant learning methods to improve generalization under distribution shift in the link prediction task. Regarding this, LPShift also demonstrates that GNN4LP as standalone models are vulnerable to distribution shift induced by pairwise heuristics. Given the severity of structural shift induced by LPShift, future work can consider relaxing the threshold between sample distributions or consider mixing structural heuristics to better mimic real-world distribution shift and the complexities therein.
Acknowledgements.
All authors are supported by the National Science Foundation (NSF) under grant numbers DGE2244164, CNS2321416, IIS2212032, IIS2212144, IOS2107215, DUE2234015, CNS2246050, DRL2405483 and IOS2035472, US Department of Commerce, Gates Foundation, the Michigan Department of Agriculture and Rural Development, Amazon, Meta, and SNAP.
Appendix A Additional Training Details
  • Please consult the project README for building the project, loading data, and re-creating results.
  • All models and link predictors are fixed to 3 layers in their neural architectures.
  • GCN and LPFormer were tuned to 128 hidden dimensions.
  • BUDDY, NCNC, NeoGNN, and SEAL were tuned to 256 hidden dimensions.
  • We apply the ’NCNC2’ variant of NCNC with an added depth argument of 2 (Wang et al., 2023a) for all CN and SP splits of ogbl-collab.
  • Otherwise, we set depth equal to 1, so as to reduce runtime.
  • NeoGNN use 1-hop neighborhoods on all tested datasets.
  • The learning rate and dropout for initial batch size tuning was fixed at $1e^{-3}$ and $0.1$ respectively.
[…]
@@ Appendix A Additional Training Details
Common Neighbors, Preferential-Attachment and Shortest-Path, as respectively shown in Figure 1, are interchangeable within the dataset splitting strategy. Details about how Common Neighbors functions within the strategy are included in Section 3.2.
@@
As another possible component of LPShift, Preferential-Attachment determines the degrees between a given source and target node and then multiples the two to produce the score, based on that score, the sample is then sorted into a new dataset split.
@@
As another possible component of LPShift, Shortest-Path determines the score by determining the minimum number of nodes necessary to reach the target node from the source node. If there is a link between the two nodes, we remove the link and then re-add to the adjacency matrix after the score calculation. The final Shortest-Path score applies the calculated shortest-path length, $SP(u,v)$ as the denominator in a ratio of $\frac{1}{SP(u,v)}$, which is then used to sort the sample into it’s respective dataset split.
Appendix B Cosine Similarity of ogbl-collab
Given LPShift’s effect on structure, we also quantify the extent to which it can affect the feature distributions of the real-world features within the ogbl-collab data. Note: ogbl-ddi was not considered due to a lack of node features and ogbl-ppa consists of one-hot encoded features derived directly from structure.
Split Train/Test Train/Valid Valid/Test
Original 83.50 ± 7.33 83.40 ± 7.33 86.91 ± 6.58
CN (0,1,2) 87.73 ± 6.12 87.42 ± 6.17 91.78 ± 4.01
(0,2,4) 87.06 ± 6.08 86.72 ± 6.14 91.01 ± 4.29
(0,3,5) 86.45 ± 6.12 86.19 ± 6.12 90.42 ± 4.41
(2,1,0) 85.53 ± 6.74 85.22 ± 6.79 90.45 ± 4.49
(4,2,0) 85.97 ± 7.09 85.55 ± 7.15 90.71 ± 4.43
(5,3,0) 86.08 ± 7.29 85.62 ± 7.34 90.74 ± 4.53
SP ($\infty$,6,4) 84.11 ± 6.94 82.12 ± 7.34 85.60 ± 6.75
($\infty$,4,3) 84.35 ± 6.93 82.66 ± 7.45 85.97 ± 6.75
(4,6$\infty$) 81.10 ± 7.74 82.15 ± 7.79 81.86 ± 7.48
(3,4$\infty$) 81.99 ± 7.84 82.81 ± 7.98 82.76 ± 7.72
PA (0,50,100) 83.92 ± 7.22 82.59 ± 7.63 85.50 ± 7.14
(0,100,200) 87.48 ± 6.16 87.55 ± 6.10 92.52 ± 3.70
(0,150,250) 87.61 ± 6.10 87.81 ± 6.17 92.45 ± 3.84
(100,50,0) 82.28 ± 7.96 83.22 ± 8.00 83.17 ± 7.85
(200,100,0) 84.75 ± 6.65 85.32 ± 6.64 89.94 ± 4.52
(250,150,0) 85.08 ± 6.65 85.51 ± 6.64 90.62 ± 4.46
Appendix C Time to Split Tested Dataset Samples
Split ogbl-collab ogbl-ppa ogbl-ddi
CN (0, 1, 2) 7.48 s 177.89 s 13.45 s
(0, 2, 4) 7.49 s 177.09 s 14.21 s
(0, 3, 5) 7.63 s 178.23 s 13.64 s
(2, 1, 0) 7.66 s 184.98 s 13.31 s
(4, 2, 0) 7.63 s 186.49 s 13.71 s
(5, 3, 0) 7.86 s 185.95 s 13.42 s
SP ($\infty$, 6, 4) 53.12 s 2748.64 s
($\infty$, 4, 3) 52.24 s 2705.91 s
(4, 6, $\infty$) 53.93 s 2715.34 s
(3, 4, $\infty$) 53.7 s 2751.56 s
PA (0, 50, 100) 19.25 s 406.5 s 15.41 s
(0, 100, 200) 19.05 s 408.04 s 14.55 s
(0, 150, 250) 19.35 s 407.81 s 13.96 s
(100, 50, 0) 19.65 s 425.3 s 14.42 s
(200, 100, 0) 19.34 s 409.93 s 14.59 s
(250, 150, 0) 19.37 s 403.55 s 14.05 s
Appendix D Size of Dataset Samples
In order to aid reproducibility of results when running LPShift manually; this section details the number of training, validation, and test edges for all LPShift dataset split types introduced within in Section 3 and all datasets.
Table 6. The number of ogbl-collab samples for the forward and backward heuristic splits.
CN (0,1,2) (0,2,4) (0,3,5) (2,1,0) (4,2,0) (5,3,0) Train 57638 237928 493790 1697336 1193456 985820 Valid 6920 20045 31676 23669 24097 25261 Test 4326 14143 21555 9048 11551 11760 SP ($\infty$,6,4) ($\infty$,4,3) (4,6$\infty$) (3,4$\infty$) Train 46880 52872 1882392 1877626 Valid 1026 1238 5222 4384 Test 2759 3457 2626 7828 PA (0,50,100) (0,100,200) (0,150,250) (100,50,0) (200,100,0) (250,150,0) Train 210465 329383 409729 1882392 1877626 457372 Valid 46626 62868 75980 64729 64202 65323 Test 9492 25527 39429 41381 30983 30999
Table 7. The number of ogbl-ppa samples for the forward and backward heuristic splits.
CN (0,1,2) (0,2,4) (0,3,5) (2,1,0) (4,2,0) (5,3,0) Train 2325936 3000000 3000000 3000000 3000000 3000000 Valid 87880 95679 98081 96765 93210 92403 Test 67176 83198 88778 92798 85448 81887 SP ($\infty$,6,4) ($\infty$,4,3) (4$\infty$,6) (3,4$\infty$) Train 17464 134728 3000000 3000000 Valid 149 4196 90511 97121 Test 20 1180 458 74068 PA (0,5k,10k) (0,10k,20k) (0,15k,25k) (10k,5k,0) (20k,10k,0) (25k,15k,0) Train 95671 98562 99352 3000000 3000000 3000000 Valid 95671 98562 99352 90623 89671 91995 Test 45251 63178 72382 44593 34321 35088
Table 8. The number of ogbl-ddi samples for the forward and backward heuristic splits.
CN (0,1,2) (0,2,4) (0,3,5) (2,1,0) (4,2,0) (5,3,0) Train 3938 5636 6580 10000 10000 10000 Valid 1184 992 942 966 1668 1462 Test 212 525 809 1808 2296 2832 PA (0,5k,10k) (0,10k,20k) (0,15k,25k) (10k,5k,0) (20k,10k,0) (25k,15k,0) Train 5132 8246 11718 10000 10000 10000 Valid 1572 3208 5058 2678 6188 6300 Test 419 1422 3694 1900 3632 5970
Appendix E Is LPShift effective at inducing distribution shift?
The following section will explore the capability of LPShift to induce a measurably-significant distribution shift in the structure of the ogbl-collab dataset. We apply the 2-sample Kolmgorov-Smirnov (KS) test (Hodges Jr, 1958) to compare if training, validation, and testing distributions can be sampled from one another, both before and after applying LPShift. As a controlled baseline to test LPShift, we generate a Holme-Kim (HK) graph (Holme and Kim, 2002) with a 40% chance to close a triangle, allowing the HK graph to contain numerous Common Neighbors without becoming fully-connected. HK graph generation parameters are included in Appendix B.